# Generalized Bayesian Inference for Causal Effects using Covariate Balancing Procedure

### v2

@@ 1 Introduction
Adjusting for covariates (or confounders) is important in estimating the causal effects in observational studies. In the case where all covariates requiring adjustment are observed, it is possible to obtain a consistent estimator for causal effects, commonly known as “no unmeasured confounding[9].” In this context, the propensity score plays a central role in estimating causal effects[25]. However, because the propensity score is usually unknown, it must be estimated using appropriate procedures, and estimators may be biased if the propensity score model is misspecified[18]. Therefore, constructing a valid model is crutial for accurate estimation.
@@
Imai and Ratkovic[16] propose the method known as the “Covariate Balancing Propensity Score” (CBPS), which does not involve modeling the propensity score directly, but, instead estimates the propensity score satisfying a moment restriction. Specifically, the moment condition comprises covariates that are desired to have a balanced distribution between the treatment groups, and the propensity score model is estimated to achieve covariate balance. The CBPS estimation heavily relies on selecting the appropriate moment restriction. Imai and Ratkovic recommend estimating CBPS based on second-order moments of covariates to achieve balance. Huang et al.[11] argue that achieving balance for higher-order moments may be necessary to obtain more accurate estimates of causal effects. More recently, to achieve a more precise covariate balance, several methods have been proposed. For example, kernel balancing methods[34, 8] aim to improve the balance by using kernel functions. Additionally, a method that satisfies the balancing of the characteristic function of the covariates has been proposed[12]. However, their methods require exact fulfillment of covariate balancing conditions, implying an equational constraint on a moment condition is necessary. Zubizarreta[36] argues that achieving an exact balance is challenging because the balance condition depends on the law of large numbers, which is applicable only in infinite samples. In other words, the condition does not account for the uncertainty derived from random sampling and treatment allocation. To address this issue, Zubizarreta proposes the “stable balancing weights” (SBW) method[36], which ensures a covariate balance up to a specified upper bound $\lambda$ that can be determined by researchers. In other words, the covariate balancing condition is subject to the subjective decision of the researcher. It is worth noting that Tan[31] demonstrates that the propensity score obtained using a logistic regression model that satisfies the same condition as the SBW method can be considered as the solution for a loss function. This result is highly relevant and critical for our proposed procedure.
@@
The SBW-based method depends on moment conditions for covariate balance and the upper bound parameter $\lambda$, which determines the acceptable range of deviation for the covariate balancing condition. As mentioned in the previous paragraph, the upper bound is subjectively determined by the researcher. This study adopts the idea of Bayesian methods to propose a methodology that relaxes the decision process of the parameter. Specifically, we leverage the conclusion that a causal effect estimator (we consider the inverse probability weighting [IPW] estimator) and the SBW-based method can be described as solutions for a loss function. Then, we proceed to specify prior distributions for each model. Additionally, to achieve the mentioned objective, we specify a prior distribution for $\lambda$. To derive the posterior distribution, which plays a central role in Bayesian inference, we employ the general Bayes procedure [3]. Using the Bayesian inference procedure, our proposed method can construct credible intervals and perform statistical tests within the Bayesian framework, not only for the causal effects of interest but also for the parameter $\lambda$.
@@
To the best of our knowledge, this study is the first to estimate the IPW estimator from a Bayesian perspective with a covariate balancing condition. The remainder of this paper is organized as follows. In Section 2, we introduces the notations and significant previous works. We specifically demonstrate that the SBW-based propensity score estimation method and IPW estimator can be explained through certain loss functions. In Section 3, we propose a Bayesian-based estimation procedure using the General Bayes paradigm. Section 4 presents simulation experiments to evaluate the performance of our proposed method against that of some competitors. In Section 5, we apply the methods to real data: Whitehall dataset.
@@ 2 Preliminaries
In this section, we introduce covariate balancing methods and their properties, specifically CBPS[16] and the SBW method[36, 31]. Furthermore, we demonstrate that the IPW estimator, used for estimating causal effects, can be considered as the solution for a loss function. Although this information is well-known, it forms a crucial foundation for the subsequent discussions in this section.
@@
In observational studies, the propensity score is usually unknown[17]. Therefore, we need to specify the model and derive the predictor of the propensity score for the subsequent analysis. Generally, the propensity score model is assumed to be the logistic regression model. However, model misspecification can lead to crucial bias in the main analysis, that is, when estimating the causal effects. To overcome the problem, Imai and Ratkovic propose CBPS, which involves estimating the parameter $\bm{\alpha}$ of the propensity score model $e(\bm{X}_{i};\bm{\alpha})$ by satisfying the following moment restriction[16]:
@@
where $g(\cdot)$ is a measurable finite dimensional ($L$ dimension) function related to the covariates; for instance,
@@
where $\lambda$ is a hyper-parameter that determines the upper bound of the covariate balance. Zubizarreta demonstrates that the upper bound $\lambda$ influences the bias of the mean of interest for the marginal expectation of outcome [36]. Theoretically, Wang and Zubizarreta show that this upper bound must converge to 00 with a certain probability order to ensure accuracy [32]. Furthermore, they discuss a bootstrap sample-based algorithm for determining $\lambda$.
@@ 2.1 Loss Function for the Propensity Score Estimation using Covariate Balancing Procedure
Note that the relationship between (2.1), and (2.2) and (2.3) is derived from the duality of a constrained optimization problem.
@@ 3 Bayesian Estimating Procedures
To construct the posterior inference procedure for $\theta_{1}$ and $\theta_{0}$, we formulate the general posterior
@@
Normalized constants $C_{1}$, $C_{0}$, and $C_{a}$ are independent of $\theta_{1}$, $\theta_{0}$, and $\bm{\alpha}$ and $\lambda$, respectively. The priors of $\theta_{1}$, $\theta_{0}$, $\bm{\alpha}$, and $\lambda$ are denoted by $p(\theta_{1}|\bm{\alpha})$, $p(\theta_{0}|\bm{\alpha})$, and $p(\bm{\alpha}|\lambda)$, and $p(\lambda)$, respectively. The $\omega$ denotes the learning rate, and the method to determine its value is described later. Note that assuming a uniform prior for $\theta_{1}$ and $\theta_{0}$ and a Laplace (or double exponential) prior with a location parameter of 0 and fixed scale parameter of $\lambda$ for $\bm{\alpha}$, the maximum a posteriori (MAP) estimator of the general posterior (3.1), where $\omega=1$ is equivalent to the estimator obtained by the estimation procedure through loss functions (2.2), (2.5), and (2.6) described in Section 2.
@@
The sampling of $\bm{\alpha}$ and $\lambda$ in Step.1 can be implemented using the weighted Bayesian bootstrap[23] or probabilistic programming language Stan[5]. In the implementation, we use the Gamma prior for $\lambda$. For Step.2, $\theta_{k}$ can be drawn from the normal distribution. Suppose that the prior of $\theta_{k}$ is the Normal distribution with the mean $\mu_{k}$ and precision $\tau_{k}$, denoted as $N(\mu_{k},\tau_{k})$. The general posterior of $\theta_{k}$ is expressed as
@@
$\begin{split}{\rm PCIC}={}&\frac{1}{n}\sum_{i=1}^{n}\frac{1}{R}\sum_{r=1}^{R}\nu(Y_{i},\bm{X}_{i}|\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\\ {}&-\left[\frac{1}{n}\sum_{i=1}^{n}\frac{1}{R}\sum_{r=1}^{R}\left\{\nu(Y_{i},\bm{X}_{i}|\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})s(Y_{i},\bm{X}_{i}|\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\right\}\right.\\ {}&\quad\quad-\left.\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{1}{R}\sum_{r=1}^{R}\nu(Y_{i},\bm{X}_{i}|\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\right\}\left\{\frac{1}{R}\sum_{r=1}^{R}s(Y_{i},\bm{X}_{i}|\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\right\}\right]\end{split}$
@@ 4 Simulation Experiments
In this section, we compare the performance of our proposed method with that of competing methods using simulation datasets. The setting is based on Leacy and Stuart[19], Setodji et al[28], and Orihara et al.[24]. The iteration number for all simulation examples was $10,000$.

### v3

@@ 1 Introduction
Adjusting for covariates (or confounders) is important in estimating the causal effects in observational studies. In the case where all covariates requiring adjustment are observed, it is possible to obtain a consistent estimator for causal effects, commonly known as “no unmeasured confounding.”[1] In this context, the propensity score plays a central role in estimating causal effects[2]. However, because the propensity score is usually unknown, it must be estimated using appropriate procedures, and estimators may be biased if the propensity score model is misspecified[3]. Therefore, constructing a valid model is crutial for accurate estimation.
@@
Imai and Ratkovic[4] propose the method known as the “Covariate Balancing Propensity Score” (CBPS), which does not involve modeling the propensity score directly, but, instead estimates the propensity score satisfying a moment restriction. Specifically, the moment condition comprises covariates that are desired to have a balanced distribution between the treatment groups, and the propensity score model is estimated to achieve covariate balance. The CBPS estimation heavily relies on selecting the appropriate moment restriction. Imai and Ratkovic recommend using the CBPS based on second-order moments of covariates to achieve balance. Huang et al.[5] argue that achieving balance for higher-order moments may be necessary to obtain more accurate estimates of causal effects. More recently, to achieve a more precise covariate balance, several methods have been proposed. For example, kernel balancing methods[6, 7] aim to improve the balance by using kernel functions. Additionally, a method that satisfies the balancing of the characteristic function of the covariates has been proposed[8]. However, their methods require exact fulfillment of covariate balancing conditions, implying an equational constraint on a moment condition is necessary. Zubizarreta[9] argues that achieving an exact balance is challenging because the balance condition depends on the law of large numbers, which is applicable only in infinite samples. In other words, the condition does not account for the uncertainty derived from random sampling and treatment allocation. To address this issue, Zubizarreta proposes the “stable balancing weights” (SBW) method[9], which ensures a covariate balance up to a specified upper bound $\lambda$ that can be determined by researchers. In other words, the covariate balancing condition is subject to the subjective decision of the researcher. It is worth noting that Tan[10] demonstrates that the propensity score obtained using a logistic regression model that satisfies the same condition as the SBW method can be considered as the solution for a loss function. This result is highly relevant and critical for our proposed procedure.
@@
The SBW-based method depends on moment conditions for covariate balance and the upper bound parameter $\lambda$, which determines the acceptable range of deviation for the covariate balancing condition. As mentioned in the previous paragraph, the upper bound is subjectively determined by the researcher. This study adopts the idea of Bayesian methods to propose a methodology that relaxes the decision process of the parameter. Specifically, we leverage the conclusion that a causal effect estimator (we consider the inverse probability weighting [IPW] estimator) and the SBW-based method can be described as solutions for a loss function. Then, we proceed to specify prior distributions for each model. Additionally, to achieve the mentioned objective, we specify a prior distribution for $\lambda$. To derive the posterior distribution, which plays a central role in Bayesian inference, we employ the general Bayes procedure [11], which updates the prior belief with loss functions. Using the Bayesian inference procedure, our proposed method can construct credible intervals and perform statistical tests within the Bayesian framework, not only for the causal effects of interest but also for the parameter $\lambda$.
@@
To the best of our knowledge, this study is the first to estimate the IPW estimator from a Bayesian perspective with a covariate balancing condition. The remainder of this paper is organized as follows. In Section 2, we introduces the notations and significant previous works. We specifically demonstrate that the SBW-based propensity score estimation method and the IPW estimator can be explained through certain loss functions. In Section 3, we propose a Bayesian-based estimation procedure using the General Bayes paradigm. Section 4 presents simulation experiments to evaluate the performance of our proposed method against that of some competitors. In Section 5, we apply the methods to real data: Whitehall dataset.
@@ 2 Preliminaries
In this section, we introduce covariate balancing methods and their properties, specifically the CBPS[4] and SBW method[9, 10]. Furthermore, we demonstrate that the IPW estimator, used for estimating causal effects, can be considered as the solution for a loss function. Although this information is well-known, it forms a crucial foundation for the subsequent discussions in this section.
@@
In observational studies, the propensity score is usually unknown[13]. Therefore, we need to specify the model and derive the predictor of the propensity score for the subsequent analysis. Generally, the propensity score model is assumed to be the logistic regression model. However, model misspecification can lead to crucial bias in the main analysis, that is, when estimating the causal effects. To overcome the problem, Imai and Ratkovic propose the CBPS, which involves estimating the parameter $\bm{\alpha}$ of the propensity score model $e(\bm{X}_{i};\bm{\alpha})$ by satisfying the following moment restriction[4]:
@@
where $g(\cdot)$ is a measurable finite dimensional ($L$ dimension) function related to the covariates; for instance (here, $L=2p$),
@@
Note that when the propensity score model is correctly specified or misspecified locally, the CBPS can derive the consistent estimator for the coefficients $\bm{\alpha}$ [Fa2022].
@@
where $\lambda$ ($>0$) is a hyper-parameter that determines the upper bound of the covariate balance. In the special case, when $\lambda=0$, the SBW-based method is obviously consistent with the CBPS. The $\lambda$ can be considered as a trade-off parameter between achieving exact balance for samples and maintaining sampling and treatment allocation randomness[9]. Zubizarreta demonstrates that the upper bound $\lambda$ influences the bias of the mean of interest for the marginal expectation of outcome [9]. Theoretically, Wang and Zubizarreta show that this upper bound must converge to 00 with a certain probability order to ensure consistency [14]. Furthermore, they discuss a bootstrap sample-based algorithm for determining $\lambda$.
@@ 2.1 Loss Function for the Propensity Score Estimation using Covariate Balancing Procedure
Note that the relationship between (2.1), and (2.2) and (2.3) is derived from the duality of a constrained optimization problem. For more details, see Appendix A.
@@ 3 Bayesian Estimating Procedures
3.1 Brief Introduction of General Bayes
First of all, we briefly introduce the general Bayesian updating proposed by Bissiri et al.[11] and connect it to the discussion in this manuscript. In this framework, to incorporate data information into prior beliefs, likelihood functions are not necessarily required; instead, a broader class of loss functions is accepted. The posterior, referred to as the generalized posterior, is defined as
$p_{G}(\theta|\bm{x})\propto p(\theta)\exp\{-\omega\ell(\theta,\bm{x})\},$
where $\bm{x}=(x_{1},\ldots,x_{n})$ is an observed sample, $p(\theta)$ is the prior distribution of the parameter $\theta$, and $\ell(\theta,\bm{x})$ is a loss function satisfying the additivity, i.e., $\ell(\theta,\bm{x})=\sum_{i=1}^{n}\ell(\theta,x_{i})$. The learning rate $\omega$, also referred to as the loss scale or calibration weight, is a non-negative scalar that regulates the learning for the parameter $\theta$ from the observed data. If $\omega$ is too large, the posterior distribution will concentrate excessively, inflating the importance of the data information about $\theta$. Conversely, if $\omega$ is too small, the posterior distribution will underweight the information from the data relative to the prior. Note that Bayes’ rule is recovered as a special case when $\omega=1$ and $\ell(\theta,\bm{x})$ is the negative log-likelihood function. Although this framework was suggested by some earlier works[Yi2009], Bissiri et al. provided a proof of the coherence and rationality of Bayesian inference, even when using general loss functions[11]. Therefore, the use of the loss functions (2.2), (2.5), and (2.6) is justified within the context of Bayesian inference.
3.2 Posterior of General Bayesian Estimation
To construct the posterior inference procedure for $\theta_{1}$ and $\theta_{0}$, we formulate the generalized posterior
@@ 3.2 Posterior of General Bayesian Estimation
Normalized constants $C_{1}$, $C_{0}$, and $C_{a}$ are independent of $\theta_{1}$, $\theta_{0}$, and $\bm{\alpha}$ and $\lambda$, respectively. The priors of $\theta_{1}$, $\theta_{0}$, $\bm{\alpha}$, and $\lambda$ are denoted by $p(\theta_{1}\mid\bm{\alpha})$, $p(\theta_{0}\mid\bm{\alpha})$, and $p(\bm{\alpha}\mid\lambda)$, and $p(\lambda)$, respectively. The $\omega$ denotes the learning rate, and the method to determine its value is described later. Note that assuming a uniform prior for $\theta_{1}$ and $\theta_{0}$ and a Laplace (or double exponential) prior with a location parameter of 0 and fixed scale parameter of $\lambda$ for $\bm{\alpha}$, the maximum a posteriori (MAP) estimator of the general posterior (3.1), where $\omega=1$ is equivalent to the estimator obtained by the estimation procedure through loss functions (2.2), (2.5), and (2.6) described in Section 2.
@@
The sampling of $\bm{\alpha}$ and $\lambda$ in Step.1 can be implemented using the loss likelihood bootstrap[15] or probabilistic programming language Stan[16]. In the implementation, we use the Gamma prior for $\lambda$.
For Step.2, $\theta_{k}$ can be drawn from the normal distribution. Suppose that the prior of $\theta_{k}$ is the Normal distribution with the mean $\mu_{k}$ and precision $\tau_{k}$, denoted as $N(\mu_{k},\tau_{k})$. The general posterior of $\theta_{k}$ is expressed as
@@
The derivation of the posterior distribution is provided in Appendix A.
In the proposed sampling schemes, the posterior sampling scheme is conducted in two stages. First, parameters related to the posterior of the propensity score model are sampled, followed by parameters for the IPW estimator. This approach corresponds to the “two-step inference” known in the Bayesian context [Zi2014, 31, 30]. Another Bayesian procedure applied in the context of causal inference is called “joint modeling”; however, these methods have both strengths and weaknesses [Li2023]. In this manuscript, we select the former procedure.
@@
$\begin{split}{\rm PCIC}={}&\frac{1}{n}\sum_{i=1}^{n}\frac{1}{R}\sum_{r=1}^{R}\nu(Y_{i},A_{i},\bm{X}_{i}\mid\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\\ {}&-\left[\frac{1}{n}\sum_{i=1}^{n}\frac{1}{R}\sum_{r=1}^{R}\left\{\nu(Y_{i},A_{i},\bm{X}_{i}\mid\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})s(Y_{i},A_{i},\bm{X}_{i}\mid\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\right\}\right.\\ {}&\quad\quad-\left.\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{1}{R}\sum_{r=1}^{R}\nu(Y_{i},A_{i},\bm{X}_{i}\mid\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\right\}\left\{\frac{1}{R}\sum_{r=1}^{R}s(Y_{i},A_{i},\bm{X}_{i}\mid\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\right\}\right]\end{split}$
@@
By applying the PCIC, the predicted risks associated with the loss functions (2.2), (2.5), and (2.6) is minimized. Focusing on the loss functions related to the causal effect ((2.5) and (2.6)), this process is similar to the one discussed in Saarela et al.[Sa2015]. Specifically, they consider a predicted (weighted) utility function for the target parameters; in their procedure, the utility is defined as a likelihood function related to the outcome. For instance, for the treatment group, the loss function becomes:
$U(\theta_{1})=\sum_{i=1}^{n}\frac{A_{i}}{e_{i}(\hat{\bm{\alpha}})}\log\left[\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left\{-\frac{1}{2\sigma^{2}}\left(Y_{i}-\theta_{1}\right)^{2}\right\}\right]\propto\sum_{i=1}^{n}\frac{A_{i}}{e_{i}(\hat{\bm{\alpha}})}\left(Y_{i}-\theta_{1}\right)^{2},$
where $\sigma^{2}$ is the variance parameter of the outcome. It can be viewed that their approach shares the same spirit as our PCIC-based procedure.
@@ 4 Simulation Experiments
In this section, we compare the performance of our proposed method with that of competing methods using simulation datasets. The setting is based on Leacy and Stuart[25], Setodji et al[26], and Orihara et al.[27]. The iteration number for all simulation examples was $10,000$. Some materials are appeared in Appendix B.